Saudi saved 180 coins, consisting of 5 - peso and 1 - peso coins. The total amount saved is 720 pesos. How many coins did each kind did he save.
Answer by Cromlix(4381)

Hi there,
Make 5 - peso coins = x
1 - peso coins = y
x + y = 180.........Eq(1)
5x + y = 720........Eq(2)
Subtract Eq(1) from Eq(2)
4x = 540
x = 135
Substitute x = 135 in Eq(1)
x + y = 180
135 + y = 180
y = 180 - 135
y = 45
Saudi has 135 5 - peso coins
and 45 1 - peso coins.
Hope this helps :-)